OpenStudy (anonymous):

Solve the equation be sure to show your work all final answers must be in simplest form. n-1/10=10

OpenStudy (hunus):

\[\frac{n-1}{10} = 10\] \[10(\frac{n-1}{10}) = 10*10\] \[n-1 = 100\] \[(n-1)+1=100+1\] \[n=101\] now lets check \[n=101\] \[\frac{101-1}{10}=10\] \[10=10\]